OpenStudy (anonymous):

Explain how the Quotient of Powers was used to simplify this expression. 2 to the fifth power, over 8 = 22

OpenStudy (zzr0ck3r):

\(\frac{2^5}{8}=\frac{2^5}{2^3}=2^{5-3}=2^2=4\)

OpenStudy (anonymous):

don't get it

OpenStudy (zzr0ck3r):

because the question is wrong \(\frac{2^5}{8}\ne 22\)